# The exact travelling wave solutions of a KPP equation

### v1

@@ I Introduction
Fisher equation admits traveling wave solutions
$u(x,t)=u(\xi),\hskip 14.22636pt\xi=x-ct,$ (3)
where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=u\left(u-1\right).$ (4)
The aim of the present paper is to show that the generalization of Eq. (4), to be presented below, can be integrated in elementary functions for the specific values of the speed $c$.
@@ II Travelling waves
We consider the equation
$u_{t}-u_{xx}=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (6)
where $n$ and $a$ are some positive constants. Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (6) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath . For the solutions of (6) in the form (3), $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right)$ (7)
with the boundary conditions (11). Equation (7) has a simple mechanical interpretation. It describes motion of the Newtonian particle with the mass 1 ($u$ being the coordinate of the particle and $\xi$ the time) in the potential well $U(u)$ defined by the equation
$\frac{dU}{du}=-u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (8)
@@
$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (10)
The boundary conditions for (7) were
$\lim_{\xi\to-\infty}u(\xi)$ $=1,$ (11a)
$\lim_{\xi\to+\infty}u(\xi)$ $=0.$ (11b)
The boundary conditions for (10) are
@@
One can easily check up that for
$c=\frac{(n+1)a+1}{\sqrt{n+1}},$ (13)
the solution of (10) satisfying the boundary conditions (12) is
$p=\frac{u\left(u^{n}-1\right)}{\sqrt{n+1}}.$ (14)
Substituting $p(u)$ into (9) and integrating we obtain the solution of (7) as
$u(\xi)=\frac{1}{\left[\exp\left(n\xi/\sqrt{n+1}\right)+1\right]^{1/n}}.$ (15)
@@
III Perturbation series
In the previous Section we used the method of integration based on Eq. (10) to find the exact integral of Eq. (7) for the specific value of $c$. Now we want to consider the limiting case $c\gg 1$ and show how the former equation can be used to construct the perturbation series for the solution of the latter.
Expanding the solution of (10) with respect to the parameter $1/c$
$p(u)=\sum_{k=1}\frac{1}{c^{k}}p^{(k)}(u),$ (17)
we obtain
$p^{(1)}$ $=u\left(u^{n}-1\right)\left(u^{n}+a\right),$ (18a)
$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u}.$ (18b)
To make our life simpler, further on consider the particular case of (10) corresponding to the Fisher equation:
$pp_{u}+cp=u(u-1).$ (19)
In this case the first two terms of the expansion (17) are
$p^{(1)}$ $=u(u-1),$ (20a)
$p^{(2)}$ $=u(u-1)(1-2u).$ (20b)
Substituting $p=p^{(1)}+p^{(2)}/c$ into (9) we obtain after a bit of algebra debnath
$u(\xi)$ $=$ $\frac{1}{\exp\left(\xi/c\right)+1}$ (21)
$+$ $\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$
IV Discussion
Let us return to (10). We have the first order ODE with two boundary conditions (12). The first impression is that the problem is overdetermined, and the solution obtained in Section II exists due to the finetuning of the l.h.s. of the equation (by choosing $c$ given by Eq. (13)). (It is irrelevant for the time being that the solution both of (10) and (7) turn out to be the elementary functions.) And the solution of (10) will not exist for any other value of $c$. However, the last statement immediately comes into contradiction with the results of the previous Section. So what about our first impression?
Notice that the boundary conditions (12) for Eq. (10) are exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$. So the boundary conditions should be taken with care. In what follows, we’ll understand that all boundary conditions are equal, but some are more equal than the others.
Let us return to Eq. (7) and consider the solution in the vicinity of $u=0$ and $u=1$. In the first case, the equation can be presented as
@@ IV Discussion
Let us start from analysing Eq. (23). The characteristic equation for the system is
$s^{2}-cs-n(a+1)=0,$ (24)
the roots being
$s_{1,2}$ $=\frac{c\pm\sqrt{c^{2}+4n(a+1)}}{2}.$ (25)
(Further on $s_{1}$ will be the positive root, and $s_{2}$ the negative.) Hence the system can be reduced to the canonical form
@@
$Y$ $=p+s_{1}v.$ (27b)
@@
Returning to Eq. (10) we realize that the boundary condition (12a) means that in the vicinity of $u=1$
$p(u)=s_{1}(u-1).$ (29)
@@
Let us start from the case of the real roots ($c>2\sqrt{a}$). In this case we can reduce the system (22) to
@@
$Y=CX^{\frac{s_{2}}{s_{1}}}.$ (34)
Because $s_{3}$ and $s_{4}$ have the same sign, $X$ approaches zero with $Y$, and all the integral curves pass through the origin, which is a node. (Note that $c$ given by (13) obviously fulfills the inequality $c>2\sqrt{a}$ for any $n>0$ and $a>0$.)
For $c<2\sqrt{a}$ the characteristic equation has two conjugate imaginary roots $\alpha+\beta i,\alpha-\beta i$, $\beta\neq 0$) we can reduce the system (22) to the form goursat
@@
If the characteristic equation has a double root $s=c/2$, the system (22) reduces to the form
@@
and the general integral is $Y=CX+X\ln X$. In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$. When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus.
Now we can resolve the (seeming) contradiction mentioned in the beginning of the present Section. Actually we have to solve ODE with a single boundary condition (12a) (presented as (29)). The second boundary condition (12b) is satisfied automatically

### v2

@@ II Travelling waves
Equation (3) admits traveling wave solutions
$u(x,t)=u(\xi),\hskip 14.22636pt\xi=x-ct,$ (4)
where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=F(u)$ (5)
with the boundary conditions
$\lim_{\xi\to-\infty}u(\xi)$ $=1,$ (6a)
$\lim_{\xi\to+\infty}u(\xi)$ $=0.$ (6b)
Equation (5) has a simple mechanical interpretation. It describes motion of the Newtonian particle with the mass 1 ($u$ being the coordinate of the particle and $\xi$ the time) in the potential well $U(u)$ defined by the equation
$\frac{dU}{du}=-F(u).$ (7)
@@
III The ODE which doesn’t contain explicitly the independent variable
@@
$pp_{u}+cp=F(u).$ (9)
The boundary conditions for (5) were (6). The boundary conditions for (9) are (however, see the next Section)
@@
Now let us be more specific and consider $F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right)$, where $n>0$ and $a\geq 0$ are some constants, so Eqs. (5) and (9) respectively take the form
$u_{\xi\xi}+cu_{\xi}$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (11a)
$pp_{u}+cp$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (11b)
Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (11a) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath . One can easily check up that there exists the solution of (11b) satisfying the boundary conditions (10)
$p=Pu\left(u^{n}-1\right),$ (12)
where
$P=\frac{1}{\sqrt{n+1}},$ (13)
provided that the speed of the wave is
$c=P+\frac{a}{P}.$ (14)
Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (11a) as
$u(\xi)=\frac{1}{\left[\exp(nP\xi)+1\right]^{1/n}}.$ (15)
@@
In Appendix A we’ll see how the results (12) - (15) can be generalized in the case when advection is taken into account in addition to diffusion.
IV The correct boundary conditions for Eq. (9)
In the previous Section we postulated two boundary conditions (10) for (9). However, the equation is of the first order. The first impression is that the problem is overdetermined, and the solution obtained there exists due to the finetuning of the l.h.s. of the equation. So one may think that we should finetune the speed $c$ for the solution of (11b) to exist, and the solution will not exist for a general value of $c$. In fact, didn’t we finetune $c$ in the previous Section?
However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (11b) for arbitrary $c\gg 1$ is presented. It looks like we have a problem. We have another problem with the boundary conditions (10) - they are exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$. In this Section we’ll solve both problems. We’ll show that one of the boundary conditions (10) should be specified, and the other can be discarded.
Consider the solution of Eq. (5) in the vicinity of $u=0$ and $u=1$. In the first case, the equation can be presented as
@@
Let us start from analysing (18). The system can be reduced to the canonical form
@@
$Y$ $=p+s_{1}v,$ (20b)
and $s_{1,2}$ are the roots of the characteristic equation
$s^{2}-cs-|F^{\prime}(1)|=0,$ (21)
that is
$s_{1,2}$ $=\frac{c\pm\sqrt{c^{2}+4|F^{\prime}(1)|}}{2}.$ (22)
(Further on $s_{1}$ will be the positive root, and $s_{2}$ the negative.)
@@
Returning to Eq. (9) we realize that the boundary condition (10a) should be specified as follows: In the vicinity of $u=1$
$p=s_{1}(u-1).$ (24)
For practical purposes we can sometimes continue to use this boundary condition in the previous form (10a), like we deed in Section III. The differential equation itself produces the correct condition (24). However, if we want, for example, to solve (9) numerically, the condition (24) should be used from the beginning.
@@
Let us start from the case of the real roots ($c^{2}>4F^{\prime}(0)$). In this case we can reduce the system (17) to
@@
$Y=CX^{\frac{s_{4}}{s_{3}}}.$ (29)
Because $s_{3}$ and $s_{4}$ have the same sign, $X$ approaches zero with $Y$, and all the integral curves pass through the origin, which is a node. (Note that $c$ given by (14) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
For $c^{2}<4F^{\prime}(0)$ the characteristic equation has two conjugate imaginary roots $\alpha+\beta i,\alpha-\beta i$, $\beta\neq 0$) we can reduce the system (17) to the form goursat
@@
If the characteristic equation has a double root $s=c/2$, which occurs for $c^{2}=4F^{\prime}(0)$, the system (17) reduces to the form
@@
and the general integral is $Y=CX+X\ln X$. In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e^{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$. When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus. Note that if for physical reasons $u$ should be strictly nonnegative, the focus solutions (and hence the value of speed $c$ for which $c^{2}\leq 4F(0)$) are unacceptable.
Now we understand that the second boundary condition (10b) can be discarded. It is satisfied automatically. And returning to the previous Section: we did finetune the speed $c$ for the solution of (11b) to be expressed in terms of elementary functions, and not for the solution to exist.
Appendix A Generalized Burgers–Huxley equation
The results obtained in Section II for the generalized KPP equation (11a) can be easily generalised to the generalized Burgers–Huxley equation cherniha ; kushner
$u_{t}-\lambda u^{n}u_{x}-u_{xx}=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (34)
where $\lambda$ is a constant. (The Burgers–Huxley equation is known in various fields of applied mathematics. For example, it describes transport processes in systems when diffusion and convection are equally important and nonlinear reaction–diffusion processes too.)
For the travelling wave solution of (34) we obtain instead of (11b) the equation
$pp_{u}+\lambda u^{n}p+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (35)
One can easily see that the solution (12) (and hence Eq. (15)) is valid in the present case also, same as the formula (14). The only difference is that Eq. (13) in this case takes the form
$P=\frac{\sqrt{\lambda^{2}+4(n+1)}-\lambda}{2(n+1)}.$ (36)
If we put $a=0$ in the formula (14), Eqs. (36) and (15) supply the solution of the generalized Huxley equation clarkson
$u_{t}-\lambda u^{n}u_{x}-u_{xx}=u^{n+1}\left(1-u^{n}\right).$ (37)
Appendix B Perturbation series
Let us return to Eq. (9). In this Appendix we’ll present the expansion of the solution of the equation (and hence of the solution of Eq. (5)) with respect to the powers of $1/c$ (meaningful if $c\gg 1$). Such expansion has the form
$p(u)=\sum_{k=1}\frac{1}{c^{k}}p^{(k)}(u),$ (38)
where
$p^{(1)}$ $=F(u),$ (39a)
$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u}.$ (39b)
Let us restrict ourselves by the next to leading order approximation with respect to the parameter $1/c$. Substituting thus truncated $p(u)$ into (8) we obtain
$\frac{du}{d\xi}=\frac{1}{c}F(u)\left(1-\frac{1}{c}\frac{dF(u)}{du}\right)$ (40)
or, equivalently (with the accepted precision),
$\frac{d\xi}{c}=\frac{du}{F(u)}+\frac{1}{c}\frac{dF(u)}{F(u)}.$ (41)
Integrating (41) we get
$\frac{\xi}{c}=\int\frac{du}{F(u)}+\frac{1}{c}\ln|F(u)|.$ (42)
To make our life simpler, further on consider only the particular case of $F(u)$ corresponding to the Fisher equation:
$F(u)=u(u-1).$ (43)
Performing integration in (42) we obtain
$\frac{\xi}{c}=\ln\left[\frac{1-u}{u}\right]+\frac{1}{c}\ln\left[u(1-u)\right].$ (44)
In the leading order with respect to $1/c$ the solution of (44) is
$u(\xi)=$ $\frac{1}{\exp\left(\xi/c\right)+1},$ (45)
hence with the accepted precision we can rewrite (44) as
$\frac{\xi}{c}-\frac{1}{c}\ln\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}=\ln\left[\frac{1-u}{u}\right].$ (46)
Exponentiating, solving with respect to $u$ and again expanding with respect to $1/c$ we obtain after a bit of algebra
$u(\xi)=\frac{1}{\exp\left(\xi/c\right)+1}$
$+\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$ (47)